OpenStudy (anonymous):

Identify which distributive property problem below would help you solve the following scenario: Your English class has 34 students. You would like to treat the class to cookies. You would like each person to have 3 cookies. Which expression would help you figure out how many cookies you would need to bake? 3(34 + 3) 3(20 − 4) 3(34 − 3) 3(30 + 4)

OpenStudy (jackellyn):

If you multiply 34*3 it tells you there should be 102 cookies made. So when you distribute it should equal the same 102. How many cookies would the each give you? 3(34) + 3(3) =? 3(20) -(3)(4) =? 3(34) -3(3)=? 3(30) +3(4)=?

OpenStudy (anonymous):

D

OpenStudy (jackellyn):

Exactly, good job